OpenStudy (anonymous):

Quadrilateral ABCD with vertices A(4, 3), B(4, -2), C(-4, -2) and D(-4, 3) is a rectangle, find the length of the diagonals. answer: sqrt 89 5 8 sqrt 39

OpenStudy (anonymous):

 For right angle triangle ADC, find distance between AD and DC \[(AD)^2 + (DC)^2 = (AC)^2\] or simply find AC and BD

OpenStudy (anonymous):

\[(8)^{2}+(5)^{2}=(9.43)^{2}\]

OpenStudy (anonymous):

@sangya21 is this correct?

OpenStudy (anonymous):

is the answer \[\sqrt{89}\]

OpenStudy (anonymous):

wow just wow thanks alot! >:(

OpenStudy (anonymous):

Yes it is. and sorry site was taking too long to reload